a ball is kicked with 16 m/s horizontal velocity and 12 m/s vertical velocity. at what velocity it will hit the ground? how much time it will be in air??
1
Expert's answer
2014-12-11T02:16:27-0500
Solution:
Given: Voy=12 m/s Vox=16 m/s
V=(Vox^2 + Voy^2)^0.5 = 20 m/s
Y = Y0 + Voy*t - g*t^2/2 ->
At moment hit the fround Y=0 Y0=0
So 0=Voy*t - g*t^2/2 - > t=2Voy/g = 2*12/9.81 = 24/9.81=2.45 s